# Integration of Large Language Models within Cognitive Architectures for Planning and Reasoning in Autonomous Robots

### v2

@@ Abstract
Symbolic reasoning systems have been used in cognitive architectures to provide inference and planning capabilities. However, defining domains and problems has proven difficult and prone to errors. Moreover, Large Language Models (LLMs) have emerged as tools to process natural language for different tasks. In this paper, we propose the use of LLMs to tackle these problems. This way, this paper proposes the integration of LLMs in the ROS 2-integrated cognitive architecture MERLIN2 for autonomous robots. Specifically, we present the design, development and deployment of how to leverage the reasoning capabilities of LLMs inside the deliberative processes of MERLIN2. As a result, the deliberative system is updated from a PDDL-based planner system to a natural language planning system. This proposal is evaluated quantitatively and qualitatively, measuring the impact of incorporating the LLMs in the cognitive architecture. Results show that a classical approach achieves better performance but the proposed solution provides an enhanced interaction through natural language.
@@ I Introduction
Therefore, using LLMs in robotics can bring several benefits and capabilities. They have been used for natural language interaction [6, 7] enabling robots to understand and generate human language, making it easier for users to communicate with and control robots with natural language. It also simplifies the knowledge retrieval [8, 9, 10] since LLMs have vast knowledge repositories, which can be used by robots to access information, answer questions, or provide explanations to users. Additionally, they can also be used in explainability and interpretability [11] for robotics, leveraging the narrative capabilities of the models to explain and interpret the logs generated by robots.
LLMs have recently achieved leading performance in tasks involving arithmetic and reasoning using the method of prompting known as chain-of-thought (CoT) [12], which encourages complex multi-step reasoning by providing step-by-step answer examples. Studies such as [13] recently achieved significant performance in such tasks using LLMs that are capable of zero-shot reasoning when it is added the instruction ”Let us consider this step by step” before each answer part included in the prompt.
@@
Our proposal focuses on reasoning, presenting the integration of LLMs in our existing cognitive architecture. The proposed approach seeks to update the current deliberative module, which is based on a symbolic planner, by one based on LLMs. Therefore, this paper analyzes the use of llama_ros, available in [18], in the MERLIN2 [19], cognitive architecture. It allows using offline LLMs inside robotics systems. Thus, its use is proposed as an alternative to classic deliberative models based on symbolic knowledge expressed in PDDL. Finally, a quantitative and qualitative review of its effect and impact on the overall decision-making system is also described in the paper.
@@
In robotics, reasoning [20] is the capability of logically and systematically processing the knowledge of the robot. There are several types of reasoning [21], for instance, practical reasoning or planning, to find the next best action and perform it; and theoretical reasoning, which aims at establishing or evaluating beliefs. The reasoning approach that is faced in this work is closer to planning type, generating sequences of actions to achieve the robot’s goals; and evaluating whether a specific goal is already achieved.
The evaluation of this proposal has been made in a human-robot interaction environment carried out in a mock-up apartment using a service robot. The reasoning faced in these missions consists of planning and evaluating whether a specific condition will be satisfied following the execution of a sequence of actions that commences from an initial state. Thus, we initially present this proposal as the implications of using LLMs in cognitive architectures. In other words, LLMs can be used in service robots although some could not see them as a formal reasoning mechanism.
The rest of the paper is organized as follows. Section II reviews the state of the art and the background. Section III depicts the material and methods of this work, that is the llama_ros tool, the cognitive architecture MERLIN2, the integration process, and the hardware and software setup. Section IV presents the evaluation of this work, composed of the experiment setup, the results and the discussions. Finally, Section V presents the conclusions and the future works.
@@ II Background and Related Work
In recent years, months, and weeks, not only the scientific community has seen remarkable progress in the field of generative AI and large language models (LLMs) but also the general public, generating a perception that it could be used everywhere. This section will review LLMs and their use for inference to current models.
@@
Deliberative and reasoning capabilities of LLMs. As LLMs have grown increasingly sophisticated and capable, their ability to engage in meaningful deliberation and planning has become a research subject. Deliberation and planning entail carefully considering various options, arguments, and perspectives before deciding and strategically organizing actions or steps to achieve a particular. Some works have attempted to use LLMs as planners. For instance, [35] uses PDDL for planning, while [36] explores using few-shot planning in embodied agents like robots.
@@
The most extended cognitive architecture category is the hybrid approach. For instance, HiMoP hybrid architecture is proposed in our previous works [42]. It is composed of a deliberative system that uses PDDL (Planning Domain Definition Language) [43] to represent the knowledge of the robot; a reactive system with a pool of state machines; and a motivational system, that contains all the robot needs. The world state of the robot is defined using the PDDL, which is the common language to carry out planning in robotics, while the state machines are used to implement the actions that the robot can perform. Alternatively, there are different technical approaches such as MOBAR [44] and CORTEX [45], where the knowledge of the robot is represented by a knowledge graph that holds symbolic and geometric information within the same structure or Ginés et al. [46, 47], where the reactive system is based on behavior trees [48]. These are the architectures that have guided the development of MERLIN2 [19], the one used in this paper.
@@ III Materials and Methods
The llama_ros tool [18], available in a public repository111https://github.com/mgonzs13/llama_ros, integrates llama.cpp [29] into ROS 2 through a suite of packages. It enables the execution of LLaMA-based models and other LLMs on devices with a low computing performance using integer quantization. This is achieved through a C/C++ implementation, facilitating model deployment across various platforms, including GPU support.
@@ III-C Integrating LLMs into MERLIN2
Incorporating LLMs into MERLIN2 aims to use their advanced reasoning abilities, transforming the existing Planning Layer. This modification involves substituting the current symbolic elements –specifically, the PDDL planner and the knowledge base, as illustrated in Figure 1 – with LLM-driven components.
@@ III-C2 Generating the World State
The structure of the knowledge items is detailed as follows: nodes are represented by the format “node is a type (properties)” and edges adhere to the schema “node relationship node (properties)”. The “(properties)” segment contains key-value pairs encapsulating the a information associated with a node or an edge, such as waypoint coordinates.
@@ IV-A Experimental Setup
The evaluation method is similar to the one depicted in MERLIN [53]. Thus, in this work, the experiments consist of greeting people in the rooms of an apartment using the RB1 robot, that is a service robot composed of a differential mobile base, with a LiDAR; and a torso, with an RGB-D camera and speakers. As a result, the robot must be able to navigate through the apartment and speak to people. Figure 5 shows the simulated apartment and the navigation map reused in this experiment. Moreover, the metrics used are the execution time, in seconds, required by the robot to complete the missions; and the traveled distance, in meters, covered by the robot during the execution of the missions.
Based on the presented missions, two experiments have been designed. Both of them are based on greeting people missions, but they differ in the number of missions. For each experiment, a randomized list of people that the robot needs to greet is created. Besides, half of the missions are canceled after 10 seconds of starting to evaluate the impact of switching missions. Hence, the first experiment comprises 6 missions, meaning greeting 6 times, of which 3 of them are canceled. The second involves 20 missions, that is, greeting 20 times, of which 10 of them are canceled.
For these experiments, the actions implemented for the experiments are the navigation and the greeting actions. The navigation action uses Nav2 [58] to move the robot between the waypoints. Thus, this action receives as arguments: the robot name, the source waypoint and the target waypoint. The greeting action uses a text-to-speak tool to greet the four persons. Its arguments are the robot’s name, the person’s name and the waypoint of the person.
On the other hand, the initial knowledge graph for the experiment is presented in Figure 6. It shows that the apartment is named GrannyHouse and has four rooms which corresponds with the marked waypoints in the map of Figure 5. The rooms are the entrance, bathroom, bedroom and living room. Besides, there are four people in different rooms, as shown in the graph. Finally, the robot RB1 is at the entrance.
@@ IV-C Discussion
The applicability of using a LLM for reasoning processes is evident. The experiment showcases a robot navigating and moving between points, similarly to the process conducted with the traditional PDDL framework.
@@ V Conclusions
It is worth discussing the PDDL performance revealed by the results of our experiments, which are superior to LLM in their different approaches. However, part of the community would prefer to interact in a more natural language manner, as is done with the LLM options which is achieved thanks to the quantized LLMs and the llama_ros tool integrated into MERLIN2. As a result, the cognitive architecture can present an improved human-robot interaction degree.
In future works, we propose to improve the performance by utilizing smaller and more accurate LLMs and specific embedding models to improve the RAG process. Besides, we will explore the use of graph algorithms to try to optimize the world state instead of using RAG as well as the use of other symbolic techniques like anthologies. Finally, we want to include the new Visual Language Models (VLM) which can be used to extract natural language text from the images captured by the robot. These ongoing efforts can contribute to the evolution and refinement of cognitive architectures incorporating cutting-edge language models.

### v3

@@ Abstract
Symbolic reasoning systems have been used in cognitive architectures to provide inference and planning capabilities. However, defining domains and problems has proven difficult and prone to errors. Moreover, Large Language Models (LLMs) have emerged as tools to process natural language for different tasks. In this paper, we propose the use of LLMs to tackle these problems. This way, this paper proposes the integration of LLMs in the ROS 2-integrated cognitive architecture MERLIN2 for autonomous robots. Specifically, we present the design, development and deployment of how to leverage the reasoning capabilities of LLMs inside the deliberative processes of MERLIN2. As a result, the deliberative system is updated from a PDDL-based planner system to a natural language planning system. This proposal is evaluated quantitatively and qualitatively, measuring the impact of incorporating the LLMs in the cognitive architecture. Results show that a classical approach achieves better performance, but the proposed solution provides an enhanced interaction through natural language.
@@ I Introduction
Therefore, using LLMs in robotics can bring several benefits and capabilities. They have been used for natural language interaction [6, 7], enabling robots to understand and generate human language, making it easier for users to communicate with and control robots with natural language. It also simplifies the knowledge retrieval [8, 9, 10] since LLMs have vast knowledge repositories, which can be used by robots to access information, answer questions, or provide explanations to users. Additionally, they can also be used in explainability and interpretability [11] for robotics, leveraging the narrative capabilities of the models to explain and interpret the logs generated by robots.
LLMs have recently achieved leading performance in tasks involving arithmetic and reasoning using the method of prompting known as chain-of-thought (CoT) [12], which encourages complex multi-step reasoning by providing step-by-step answer examples. Studies such as [13] recently achieved significant performance in such tasks using LLMs that are capable of zero-shot reasoning when it is added the instruction Let us consider this step by step in the prompt.
@@
Our proposal focuses on reasoning and presenting the integration of LLMs in our existing cognitive architecture. The proposed approach seeks to update the current deliberative module, which is based on a symbolic planner, by one based on LLMs. Therefore, this paper analyzes the use of llama_ros, available in [18, 19], in the MERLIN2 [20, 21], cognitive architecture. It allows using offline LLMs inside robotics systems. Thus, its use is proposed as an alternative to classic deliberative models based on symbolic knowledge expressed in PDDL. Finally, a quantitative and qualitative review of its effect and impact on the overall decision-making system is also described in the paper.
@@
In robotics, reasoning [22] is the capability of logically and systematically processing the knowledge of the robot. There are several types of reasoning [23], for instance, practical reasoning or planning, to find the next best action and perform it; and theoretical reasoning, which aims at establishing or evaluating beliefs. The reasoning approach that is faced in this work is closer to a planning type, generating sequences of actions to achieve the robot’s goals; and evaluating whether a specific goal is already achieved.
The evaluation of this proposal has been made in a human-robot interaction environment carried out in a mock-up apartment using a service robot. The reasoning faced in these missions consists of planning and evaluating whether a specific condition will be satisfied following the execution of a sequence of actions that commences from an initial state. Thus, we initially present this proposal as the implications of using LLMs in cognitive architectures. In other words, LLMs can be used in service robots, although some could not see them as a formal reasoning mechanism.
The rest of the paper is organized as follows. Section II reviews the state of the art and the background. Section III depicts the material and methods of this work, that is, the llama_ros tool, the cognitive architecture MERLIN2, the integration process, and the hardware and software setup. Section IV presents the evaluation of this work, composed of the experiment setup, the results and the discussions. Finally, Section V presents the conclusions and the future work.
@@ II Background and Related Work
In recent years, months, and weeks, not only the scientific community has seen remarkable progress in the field of generative AI and large language models (LLMs) but also the general public, generating a perception that it could be used everywhere. This section will review LLMs and their use for inference in current models.
@@
Deliberative and reasoning capabilities of LLMs. As LLMs have grown increasingly sophisticated and capable, their ability to engage in meaningful deliberation and planning has become a research subject. Deliberation and planning entail carefully considering various options, arguments, and perspectives before deciding and strategically organizing actions or steps to achieve a particular goal. Some works have attempted to use LLMs as planners. For instance, [37] uses PDDL for planning, while [38] explores using few-shot planning in embodied agents like robots.
@@
The most extended cognitive architecture category is the hybrid approach. For instance, HiMoP hybrid architecture is proposed in our previous works [44]. It is composed of a deliberative system that uses PDDL (Planning Domain Definition Language) [45] to represent the knowledge of the robot; a reactive system with a pool of state machines; and a motivational system, that contains all the robot needs. The world state of the robot is defined using the PDDL, which is the common language to carry out planning in robotics, while the state machines are used to implement the actions that the robot can perform. Alternatively, there are different technical approaches such as MOBAR [46] and CORTEX [47], where the knowledge of the robot is represented by a knowledge graph that holds symbolic and geometric information within the same structure or Ginés et al. [48, 49], where the reactive system is based on behavior trees [50]. These are the architectures that have guided the development of MERLIN2 [20, 21], the one used in this paper.
@@ III Materials and Methods
The llama_ros tool [18], available in a public repository111https://github.com/mgonzs13/llama_ros, integrates llama.cpp [31] into ROS 2 through a suite of packages. It enables the execution of LLaMA-based models and other LLMs on devices with low computing performance using integer quantization. This is achieved through a C/C++ implementation, facilitating model deployment across various platforms, including GPU support.
@@ III-C Integrating LLMs into MERLIN2
Incorporating LLMs into MERLIN2 aims to use their advanced reasoning abilities, transforming the existing Planning Layer. This modification involves substituting the current symbolic elements, specifically the PDDL planner and the knowledge base, as illustrated in Figure 1, with LLM-driven components.
@@ III-C2 Generating the World State
The structure of the knowledge items is detailed as follows: nodes are represented by the format “node is a type (properties)” and edges adhere to the schema “node relationship node (properties)”. The “(properties)” segment contains key-value pairs encapsulating the information associated with a node or an edge, such as waypoint coordinates.
@@ IV-A Experimental Setup
The evaluation method is similar to the one depicted in MERLIN [59]. Thus, in this work, the experiments consist of greeting people in the rooms of an apartment using the RB1 robot, which is a service robot composed of a differential mobile base, with a LiDAR, and a torso, with an RGB-D camera and speakers. As a result, the robot must be able to navigate through the apartment and speak to people. Figure 5 shows the simulated apartment and the navigation map reused in this experiment. Moreover, the metrics used are the execution time, in seconds, required by the robot to complete the missions; and the traveled distance, in meters, covered by the robot during the execution of the missions.
Based on the presented missions, two experiments have been designed. Both of them are based on greeting people missions, but they differ in the number of missions. For each experiment, a randomized list of people that the robot needs to greet is created. Besides, half of the missions are canceled after 10 seconds of starting to evaluate the impact of switching missions. Hence, the first experiment comprises 6 missions, meaning greeting 6 times, of which 3 of them are canceled. The second involves 20 missions, that is, greeting 20 times, of which 10 are canceled.
For these experiments, the actions implemented for the experiments are the navigation and greeting actions. The navigation action uses Nav2 [60] to move the robot between the waypoints. Thus, this action receives as arguments: the robot name, the source waypoint and the target waypoint. The greeting action uses a text-to-speak tool to greet the four people. Its arguments are the robot’s name, the person’s name and the waypoint of the person.
On the other hand, the initial knowledge graph for the experiment is presented in Figure 6. It shows that the apartment is named GrannyHouse and has four rooms, which corresponds with the marked waypoints in the map of Figure 5. The rooms are the entrance, bathroom, bedroom and living room. Besides, there are four people in different rooms, as shown in the graph. Finally, the robot RB1 is at the entrance.
@@ IV-C Discussion
The applicability of using an LLM for reasoning processes is evident. The experiment showcases a robot navigating and moving between points, similarly to the process conducted with the traditional PDDL framework.
@@ V Conclusions
It is worth discussing the PDDL performance revealed by the results of our experiments, which is superior to LLM in its different approaches. However, part of the community would prefer to interact in a more natural language manner, as is done with the LLM options, which is achieved thanks to the quantized LLMs and the llama_ros tool integrated into MERLIN2. As a result, the cognitive architecture can present an improved human-robot interaction degree.
In future works, we propose to improve the performance by utilizing smaller and more accurate LLMs and specific embedding models to improve the RAG process. Besides, we will explore the use of graph algorithms to try to optimize the world state instead of using RAG, as well as the use of other symbolic techniques like anthologies. Finally, we want to include the new Visual Language Models (VLM), which can be used to extract natural language text from the images captured by the robot. These ongoing efforts can contribute to the evolution and refinement of cognitive architectures, incorporating cutting-edge language models.